OpenStudy (samsan9):

rewrite using radicals (a^3/4)^5

OpenStudy (samsan9):

\[a(\frac{ 3 }{ 4 })^5\]

OpenStudy (astrophysics):

\[\huge (a ^{\frac{ 3 }{ 4 }})^{5}?\]

OpenStudy (astrophysics):

\[\huge x ^{\frac{ m }{ n }} = \sqrt[n]{x ^{m}}\] Multiply the exponents then apply this rule.

OpenStudy (samsan9):

cool thanks you just lliterally helped me solve 2 questions lol

OpenStudy (astrophysics):

Np :)

OpenStudy (samsan9):

is it the same xm/n for the other way around?

OpenStudy (astrophysics):

Yes :) hence the equal sign ^.^